OpenStudy (anonymous):

can you please help me i really need it again. Judah made pottery to sell at his school's craft bazaar. He made $5 for each piece of pottery that he sold. Write a function rule to define the total amount y Judah made based on how many pieces x he sold. Give an example of how the function rule works. Which is the dependent variable? The independent variable?

OpenStudy (across):

Judas betrayed Jesus for 5 rupees. That means that the equation he used to sell Jesuses was something like\[\text{Rupees}=5\times\text{Jesus}.\]By the powers of Tutankhamen, you can summon the Lord Crypt to hang you on a tree upside down and tell you that the independents will fight those who still live under mommy for worthless pieces of compressed sand.

OpenStudy (anonymous):

umm i dont understand that

OpenStudy (across):

أنا لا أفهم أيضا.

OpenStudy (anonymous):

ummm and i dont understand that

OpenStudy (phi):

He made $5 for each piece of pottery that he sold. Write a function rule let n= # of pieces sold let f(n) = dollars he got after selling n pieces f(n)= 5n Here is an example of how to use it: He sold 10 pieces. How much money did he get? f(10)= 5*10= 50 The independent variable is n the dependent is f(n) which you could give a name for example, y